A plane flying 30.0 m above the sea at a velocity of 100.0 m/s along the horizontal will drop relief goods to people who are stranded in an island. When shall the goods be released so that they fall exactly where the people are located? Prove your answer through your solution.


1
Expert's answer
2021-11-29T15:13:51-0500

"v_{air }= 100\frac{m}{s}"

"h=30m"

"g = 9.8\frac{m}{s^2}"

"v_x =v_{air} =100\frac{m}{s}"

"s = v_xt"

"v_{y_0}=0"

"t =\sqrt{ \frac{2h}{g}}"

"s = v_xt= v_x\sqrt{ \frac{2h}{g}}=100\sqrt{ \frac{2*30}{9.8}}=247.44m"

"s - \text{distance traveled by the cargo horizontally before landing}"

"\text{the cargo released at a given distance from the island will land on it.}"


"\text{Answer: The cargo must be released before reaching the island } 247.44m"